# PSZ: Enhancing the SZ Scientific Lossy Compressor With Progressive Data Retrieval

### v1

@@ 4.2.2. Prediction models
Prediction models are more complex than transform models, as they tightly coupled the decorrelation function $P$ with quantization $Q$. For transform models, the whole dataset is first transformed by $T$ before applying quantization $Q$. However, in prediction models, the data usually is split into orthogonal layers or groups, and the prediction $P$ and quantization $Q$ are applied layer by layer (Di et al., 2024; Ainsworth et al., 2018), such that the input for prediction is $\hat{x}$ (data with lossy error), instead of $x$ (original data). As a result, the prediction model can be viewed as a non-linear mapping between input and decorrelate data.
As demonstrated in Figure 2, the whole linear space is divided into multiple orthogonal levels by setting shrinking stride distance. As a denote, we take the sub linear spaces $V_{l}$ of the linear space $\mathbb{R}^{n}$ in which the original dataset $x$ resides as a level. The total number of levels is defined as $L$.
$V_{l}\subseteq\mathbb{R}^{n},l=1,2,...,L$
The prediction will be executed $L-1$ times, covering from level $L$ to level 1.
$P_{l}:V_{l+1}\to V_{l},l=1,2,...,L-1$
In prediction models, we store the prediction difference $y_{l}$ in $l$-th level as quantized values $q_{l}$, instead of storing transformed values as transform models. For example, suppose we are at level $l$, $y_{l}$ is the diff between the original value $x_{l}$ and the prediction based on $\hat{x}_{l+1}$, the previous level’s data with lossy error.
@@
The diff will be quantized to an integer while ensuring the point-wise lossy error is less than the pre-defined global error bound $eb$:
$y_{l}=\hat{y}_{l}+\varepsilon_{l},\|\varepsilon_{l}\|_{\infty}\leq eb$
Then we can retrieve the next level by
@@
By simplifying the three formulas above, we can get the conclusion that for each level, $\hat{x}_{l}$ only differs with original data $x_{l}$ within the point-wise error bound $eb$
@@
4.3. Progressive interpolation solution
@@
The foundation of our progressive design is to load a partial of the information of $\hat{y}$ according to the bitplane. As illustrated in Figure 4, in our approach, the quantized data (denoted as $q_{l}$ at level $l$) are 32-bit integers. The bits from the same position across multiple quantized integers form a bitplane (represented by small rectangular boxes in the figure). We independently encode the 32 bitplanes within the same level, such that the output in a lower fidelity can be reconstructed by loading some, not all, of the bits. When retrieving data, our optimizer discussed in Section 5 will determine the optimal loading strategy – the minimum number of bitplanes to load from each level to meet the retrieval requirement.
@@ 4.4.1. Predictive Bitplane Coding
We propose a predict-based strategy to exploit the correlation between bitgroups. To support progressive retrieval, every integer of $\hat{y}$ is split by bitplanes, which means the bits from a single integer are encoded separately. As a result, the correlation between bits from the same integer is totally ignored. We observe that during decompression when retrieving a certain bitplane at a specific level, the previously loaded bitplanes are already known. This observation allows us to leverage the correlation between bitplanes by predicting the bit value based on previous bits from the same integer, and encode the results of the prediction instead of raw bits. For example, if the prefix bit is 0, we predict that the current bit $b$ will also be 0. If the prefix bit is 1, we predict that $b$ will be 1. This prediction mechanism is equivalent to applying an XOR operation between the prefix bit and $b$ as $\text{EncodedBit}=\text{PrefixBit}\oplus b$.
@@
$\text{EncodedBit}=(\text{PrefixBit}_{1}\oplus\text{PrefixBit}_{2}\oplus...\oplus\text{PrefixBit}_{n})\oplus b$
@@ 4.4.2. Negabinary coding
Unlike their non-progressive counterparts, progressive compression schemes must efficiently handle sign bits, since loading them is a prerequisite for reconstructing values, but storing sign bits separately and loading them first before all other bits incurs significant overhead. To address this issue, we evaluate the suitability of three widely used methods for encoding sign bits in a progressive setting – two’s complement (Patterson and Hennessy, 2013), sign-magnitude (Patterson and Hennessy, 2013), and negabinary (Hoang et al., 2018; Lindstrom, 2014), and we select negabinary encoding as the preferred approach.
First, negabinary encoding may lead to better compressibility than other methods for values that fluctuate around zero (which is predominant in the last layer of quantized integers). In two’s complement, the higher-order bits often contain many 1s for values near zero, making the corresponding high-order bitplanes difficult to compress. For example, the 8-bit representations of 1 and -1 are 00000001 and $11$ in two’s complement, 00000001 and 10000001 in sign-magnitude, and 00000001 and 00000011 in negabinary. As a result, negabinary encoding keeps the higher-order bits as 0, which leads to more compressible high-order bitplanes compared to the dense 1s in two’s complement. Second, compared to sign-magnitude encoding, negabinary encoding is more balanced. When the least significant bits (LSBs) are set to zero (due to quantization or truncation), the resulting error uncertainty is smaller in negabinary encoding. This leads to more predictable error behavior, which is beneficial for data compression and reconstruction. The uncertainty could be formulated by
@@
where n represents the number of bits discarded. As n increases, the uncertainty of negabinary is only around 2 thirds of sign-magnitude encoding.
@@ 5. Optimized data loading
In this section, we present our optimized data load strategy that can minimize the volume of data loaded while satisfying the retrieval constraints. Our optimizer supports both error-bound mode and bitrate mode which represents most of the scenarios in scientific data retrieval.
  • Error bound mode: in this mode, the user specifies the required precision and the optimizer selects the optimal strategy to ensure that the reconstructed data’s error remains strictly within the given bound while minimizing the amount of data loaded.
  • Fixed rate/size mode: in this mode, the user specifies a maximum allowable bitrate or size, typically constrained by I/O bandwidth. The optimizer then ensures that the reconstructed data has the smallest possible error while staying within the specified bitrate limit. Since the total data size is fixed, requiring a specific bitrate is equivalent to requiring a fixed retrieval size.
We first discuss the upper bound theory which both of modes rely on, then show how the optimizer make best decisions on bitplane selection while balancing accuracy and efficiency.
@@ 5.1. Fundamental theory for our optimizer
In this section, we discuss the fundamental theory for our optimizer which measures the lossy error in each layer.
@@
The $L_{\infty}$ error of our compressor is bounded by
(5) $\|x-\hat{x}\|_{\infty}\leq\sum\|P_{1}\|_{\infty}\|P_{2}\|_{\infty}...\|P_{l}\|_{\infty}\|\delta y_{l+1}\|_{\infty}+eb$
The $L_{\infty}$ of the prediction $P$ depends on the specific prediction function. For linear interpolation, its value is 1, and for cubic interpolation, it equals 1.25. As a result,  Equation 5 can also be expressed as
@@
where each coefficient $p_{l}$ could be 1 or 1.25.
@@
We start from the top level L where the progressive decompression begins. The loss or error in this level caused by loading partial bitplanes is defined as $\delta y_{L}$. The relation between $\hat{y}_{L}$, which is calculated with all the bitplanes, and the dequantized $\hat{y}_{l}^{\prime}$ is:
$\delta y_{L}=\hat{y}_{L}-\hat{y}_{L}^{\prime}$
The decompressed value at level $L$ is $\hat{x}^{\prime}_{L}=P_{L}(\mathbf{0})+\hat{y}^{\prime}_{L}$. Comparing the decompressed data with all bitplanes $\hat{x}_{L}$, we can get:
$\hat{x}_{L}=\hat{x}_{L}^{\prime}+\delta y_{L}$
As Algorithm 1 shows, in our approach, we load part of the bitplanes of each level, dequantizing it, getting $\hat{y}^{\prime}_{l}=Q^{-1}q^{\prime}_{l}$, then move on to the next level. A major challenge in such a design is that the uncertainty would propagate between levels. For example, if we move on to level $L-1$, the dequantized residual is $\hat{y}_{L-1}^{\prime}=\hat{y}_{L-1}-\delta y_{L-1}$, and if we evaluate the decompressed value $\hat{x}^{\prime}_{L-1}$ and compare it with $\hat{x}_{L-1}$, we can get:
$\hat{x}^{\prime}_{L-1}=P_{L-1}(\hat{x}_{L}-\delta y_{L})+\hat{y}_{L-1}-\delta y_{L-1}\\ =\hat{x}_{L-1}-\delta y_{L-1}-P_{L-1}\delta y_{L}$
This equation tells us the error comes with two terms. The first one $\delta y_{L-1}$ is the uncertainty in level $L-1$, and the second term $P_{L-1}\delta y_{L}$ is the error propagated from the previous level.
Applying the same analysis to other levels, we can get:
$x_{l}$ $=\hat{x}_{l}^{\prime}+\delta y_{prop,l}+\delta y_{l}+\varepsilon_{l}$
@@
Such that:
$x_{l}-\hat{x}^{\prime}_{l}=\delta y_{l}+\sum_{m=0}P_{l}P_{l+1}...P_{l+m}\delta y_{l+m+1}+\varepsilon_{l}$
In our approach, the $L_{\infty}$ of interpolation $P_{l}$ is greater or equal to 1, which implies we should expect the lower level to have larger $L_{\infty}$ errors. In other words, if we can bound the lowest level, then we can bound the error of the entire dataset. As a result, we have:
$\|x-\hat{x}\|_{\infty}$ $\leq\|x-\hat{x}^{\prime}_{1}\|_{\infty}$
@@
In terms of $L_{\infty}(P)$, its value depends on the prediction method. Liner interpolation, as shown in Equation 1, has two coefficients both at 0.5. Such that $L_{\infty}(P)=0.5+0.5=1$. Similarly, cubic interpolation, as shown in Equation 2, has four coefficients, and $L_{\infty}(P)=2\times\frac{1}{16}+2\times\frac{9}{16}=1.25$. With $L_{\infty}(P)$ being 1 or 1.25, we can simplify Equation 8 to Equation 6.
@@ 5.2. Optimized loading based on error-bounds
The goal of error bound mode is to generate a loading strategy specifying the minimum number of bitplanes to load in each layer, in order to let all point-wise lossy error to be less or equal to the user-defined bound (denoted as E). Finding such bitplanes can be formalized as a optimization problem.
@@
$b_{l}$ indicates the number of bitplanes discarded at the level $l$. $SavedSize(l,b_{l})$ denotes the amount of data saved by discarding $b_{l}$ bitplanes at level $l$. $err(l,b_{l})$ denotes the error at level $l$ by discarding $b_{l}$ bitplanes. It can be formulated by Equation 6 as $err(l,b_{l})=p^{l-1}\|\delta y_{l}\|_{\infty}$, where $p=1$ for linear interpolation or $1.25$ for cubic interpolation. $\delta y_{l}$ is the uncertainty derived from Section 4.4.2 which is a function of $b_{l}$.
@@ 5.3. Optimized loading based on bitrates
Similar with Section 5.2, we effectively convert this problem to the classic knapsack dynamic programming problem, so that we can solve it the same way as in the error bound mode, also with minimal overhead.
@@ 6.1.2. Datasets
The experiments are evaluated on six scientific datasets listed in Table 3.
@@ 6.1.3. State-of-the-Art lossy compressors in our evaluation
The experiments include four SOTA scientific lossy compressors as baseline – SZ3, SZ3-R, ZFP-R, and PMGARD.
SZ3-M (Wu et al., 2024): SZ3-M (where ”M” stands for multi fidelity) is the straightforward multi fidelity version of SZ3 based on multiple outputs. SZ3 is the leading non-progressive scientific lossy compressor. SZ3 uses interpolation as prediction, together with linear-scale quantization, Huffman coding, and zstd lossless coding. SZ3 has excellent compression ratio and fidelity over others while its speed may not be as fast as ZFP. SZ3-M compresses the input with different error bounds independently and groups those compressed data all together as output. Such a solution supports multi-fidelity retrieval but is not progressive yet, as it cannot reuse the low-fidelity data to build high-fidelity results.
@@
ZFP-R (Magri and Lindstrom, 2024): ZFP-R is the progressive version of ZFP based on residuals. It handles residuals the same way as SZ3-R. ZFP is the leading transform-based lossy compressor based on orthogonal transformation. ZFP (Lindstrom, 2014) is usually the fastest scientific lossy compressor because of its highly efficient transform function, although its compression ratio may not be as high as others.
@@
We note that the residual-based approaches (SZ3-R and ZFP-R) has drawbacks that affect our evaluation. First, they have limited error-bound flexibility. More specifically, the retrieval is only possible at a few predefined error bounds. This creates a trade-off – setting too many error bounds reduces the overall compression ratio and significantly degrades compression/decompression throughput, while setting too few nodes limits flexibility in selecting error bounds during decompression. For the experiments, we configure five error bounds for them, where each successive one is a factor of $2^{2}$ (4×) apart. Specifically, the error bounds are set to $2^{16}eb$, $2^{14}eb$ $2^{12}eb$, $2^{10}eb$ $2^{8}eb$, $2^{6}eb$, $2^{4}eb$, $2^{2}eb$, and $eb$. Second, those methods are not well-suited for retrievals based on pre-defined bitrates, as the residuals are compressed by error bounds, and their size is not aligned with the bitrate of the retrieval requests. As a result, we select the largest residual that fits within the user-specified bitrate constraint for SZ3-R and ZFP-R in the experiments.

### v2

@@ 4.2.2. Prediction models
Prediction models are more complex than transform models, as they tightly couple the decorrelation function $P$ with quantization $Q$. For transform models, the whole dataset is first transformed by $T$ before applying quantization $Q$. However, in prediction models, the data usually is split into orthogonal layers or groups, and the prediction $P$ and quantization $Q$ are applied layer by layer (Di et al., 2024; Ainsworth et al., 2018) on $\hat{x}$ (data with lossy error) instead of $x$ (original data). As a result, the prediction model can be viewed as a non-linear mapping between input and decorrelate data.
Figure 3 shows how to divide the whole linear space into multiple orthogonal levels by setting a shrinking stride distance. We represent level as the sublinear spaces $V_{l}\subseteq\mathbb{R}^{n}$. The total number of levels is defined as $L$. The prediction $P_{l}$ will be executed $L-1$ times, covering from the top level $L$ to level 1.
For each level $l$, we process $y_{l}$ which is the difference between the original data $x_{l}$ and its prediction. The prediction $P_{l}$ is based on interpolation from $\hat{x}_{l+1}$ (the previous level’s data with lossy error).
@@
The prediction difference $y_{l}$ will be quantized to integers, and dequantized back to floating point $\hat{y}_{l}$. The quantization guarantees that the point-wise error between $y$ and $\hat{y}_{l}$ is less than the pre-defined global error bound $eb$. Note that in decompression we do not have the original prediction difference $y$ since it is lossy quantized here.
$y_{l}-\hat{y}_{l}=\varepsilon_{l},\|\varepsilon_{l}\|_{\infty}\leq eb$
In decompression, which is the reverse process, we first recover the prediction data $P_{l}\hat{x}_{l+1}$, then compute the output $\hat{x}$ by adding $\hat{y}$ (the prediction difference with error) on top of the prediction.
@@
By simplifying the three formulas above, we can get the conclusion that for each level, $\hat{x}_{l}$ only differs with original data $x_{l}$ within the point-wise error bound $eb$.
@@
4.3. Progressive interpolation algorithm
@@
The foundation of our progressive design is to load a partial of the information of $\hat{y}$ according to the bitplane. As illustrated in Figure 4, in our approach, the quantized data (denoted as $q_{l}$ at level $l$) are 32-bit integers. The bits from the same position across multiple quantized integers form a bitplane (represented by small rectangular boxes in the figure). We independently encode the 32 bitplanes at the same level, which allows lower-fidelity outputs to be reconstructed by loading only some of the bits rather than all of them. When retrieving data, our optimizer discussed in Section 5 will determine the optimal loading strategy – the minimum bitplanes to load from each level to meet the retrieval requirement.
@@ 4.4.1. Predictive Bitplane Coding
We propose a predict-based strategy to exploit the correlation between bitgroups. To support progressive retrieval, each quantized integer is split into bitplanes, which means that the bits of a single integer are encoded separately. As a result, the correlation between bits from the same integer is totally ignored. We observe that during decompression when retrieving a certain bitplane at a specific level, the previously loaded bitplanes are already known. This observation allows us to leverage the correlation between bitplanes by predicting the bit value based on previous bits from the same integer, and encode the results of the prediction instead of raw bits. For example, if the prefix bit is 0, we predict that the current bit $b$ will also be 0. This prediction mechanism is equivalent to applying an XOR operation between the prefix bit and $b$ as encoded $\text{encoded bit}=\text{prefix bit}\oplus b$.
@@
[…]
@@ 4.4.2. Negabinary coding
Unlike their non-progressive counterparts, progressive compression schemes must efficiently handle sign bits, since loading them is a prerequisite for reconstructing values, but storing sign bits separately and loading them first before all other bits incurs significant overhead. To address this issue, we evaluate the suitability of three widely used methods for encoding sign bits in a progressive setting – two’s complement (Patterson and Hennessy, 2013), sign-magnitude (Patterson and Hennessy, 2013), and negabinary (Hoang et al., 2018; Lindstrom, 2014), and select negabinary encoding as the preferred approach for two reasons.
First, negabinary encoding may lead to better compressibility than other methods for values that fluctuate around zero (which is predominant in the last layer of quantized integers). In two’s complement, the higher-order bits often contain many 1s for values near zero, making the corresponding high-order bitplanes difficult to compress. For example, the 8-bit representations of 1 and -1 are 00000001 and $11$ in two’s complement, 00000001 and 10000001 in sign-magnitude, and 00000001 and 00000011 in negabinary. As a result, negabinary encoding keeps the higher-order bits as 0, which leads to more compressible high-order bitplanes compared to the dense 1s in two’s complement. Second, compared to sign-magnitude encoding, negabinary encoding is more balanced. When the least significant bits are set to zero (due to quantization or truncation), the resulting error uncertainty is smaller in negabinary encoding. This leads to more predictable error behavior, which is beneficial for data compression and reconstruction. The uncertainty could be formulated by
@@
where d represents the number of bits discarded. As d increases, the uncertainty of negabinary is only around two-thirds of sign-magnitude encoding.
@@ 5. Optimized data loading
In this section, we present our optimized data load strategy that can minimize the volume of data loaded under the retrieval constraints. Our optimizer supports both error-bound mode and bitrate mode which represents most of the scenarios in scientific data retrieval.
  • Error bound mode: in this mode, users specify the required precision and the optimizer selects the optimal strategy to ensure that the reconstructed data’s error remains strictly within the given bound while minimizing the amount of data loaded.
  • Fixed rate/size mode: in this mode, the user specifies a maximum allowable bitrate or size, typically constrained by I/O bandwidth. The optimizer then ensures that the reconstructed data has the smallest possible error while staying within the specified bitrate limit. Since the total data size $n$ is constant for the input, requiring a specific bitrate is equivalent to requiring a fixed retrieval size.
We first discuss the upper bound theory which both modes rely on, then show how the optimizer makes the best decisions on bitplane selection based on dynamic programming.
@@ 5.1. Fundamental theory for our optimizer
In this section, we discuss the fundamental theory for our optimizer which measures the error caused by only a partial of bitplanes being loaded.
@@
The $L_{\infty}$ error in progressive retrieval can be bounded based on the information loss due to the unloaded bitplanes.
@@
In this equation, $p=1$ for linear interpolation, and $p=1.25$ for cubic interpolation. $\delta y_{l}$ describes the information loss in level $l$ caused by the unloaded bitplanes, and its value can be pre-computed during compression.
@@
As Algorithm 1 shows, the retrieval process is done iteratively on all levels, with each level taking the previous level’s output for the prediction of the current level. Consequently, the major challenge here for error analysis is that the uncertainty would propagate between levels. Towards this, we model the error in level $i$ comes with three terms. The first one $\delta y_{l}$ is the information loss due to unloaded bitplanes in this level. The second term $\delta y_{prop,l}$ is the information loss propagated from the previous level. The third term $\varepsilon_{l}$ is the error caused by the lossy quantization.
(6) $x_{l}-\hat{x}_{l}=\delta y_{l}+\delta y_{prop,l}+\varepsilon_{l}$
The information loss propagated until the previous level $l+1$ is $\delta y_{prop,l+1}$. It will be first combined with the information loss in the current level $\delta y_{l+1}$, and then transfered by prediction and passed to the next level as $\delta y_{prop,l}$.
@@
Combine the two equations, we can have:
(8) $x_{l}-\hat{x}_{l}=\delta y_{l}+\sum_{m=0}P_{l}P_{l+1}...P_{l+m}\delta y_{l+m+1}+\varepsilon_{l}$
In our approach, the $L_{\infty}$ of interpolation $P_{l}$ is greater or equal to 1, which implies the information loss would be amplified to the lower level. In other words, lower levels have larger errors, and the maximum error of the dataset is in the first level. As a result, we can quantify the error of the dataset with Equation 8 by:
$\|x-\hat{x}\|_{\infty}$ $=\|x-\hat{x}_{1}\|_{\infty}$
@@
Now we simplify Equation 9 to Equation 5. The value of $L_{\infty}(P)$ depends on the prediction method. Liner interpolation, as shown in Equation 1, has two coefficients both at 0.5. Such that $L_{\infty}(P)=0.5+0.5=1$. Similarly, cubic interpolation, as shown in Equation 2, has four coefficients, and $L_{\infty}(P)=2\times\frac{1}{16}+2\times\frac{9}{16}=1.25$. With the derived value of $L_{\infty}(P)$, Equation 9 is converted to Equation 5.
@@ 5.2. Optimized loading based on error-bounds
The goal of error-bound mode is to generate a loading strategy specifying the minimum number of bitplanes to load in each layer, in order to let all point-wise lossy errors be less or equal to the user-defined bound (denoted as E). Finding such bitplanes can be formalized as an optimization problem.
@@
$b_{l}$ indicates the number of bitplanes discarded at the level $l$. $SavedSize(l,b_{l})$ denotes the amount of data saved by discarding $b_{l}$ bitplanes at level $l$. $err(l,b_{l})$ denotes the error at level $l$ by discarding $b_{l}$ bitplanes. As stated in Section 4.4.2, $err(l,b_{l})=p^{l-1}\|\delta y_{l}\|_{\infty}$, where $p=1$ for linear interpolation, $p=1.25$ for cubic interpolation, and $\delta y_{l}$ is a function of $b_{l}$.
@@ 5.3. Optimized loading based on bitrates
Similar with Section 5.2, we effectively convert this problem to the knapsack dynamic programming problem, so that it can be solved similarly as in the error bound mode, also with negligible overhead.
@@ 6.1.2. Datasets
The experiments are evaluated on six datasets across four diverse scientific domains, as listed in Table 3.
@@ 6.1.3. State-of-the-Art lossy compressors in our evaluation
The experiments include four SOTA scientific lossy compressors as baselines – SZ3-M, SZ3-R, ZFP-R, and PMGARD.
SZ3-M (Wu et al., 2024): SZ3-M (where ”M” stands for multi-fidelity) is the straightforward multi fidelity version of SZ3 based on multiple outputs. SZ3 is the leading non-progressive scientific lossy compressor. SZ3 uses interpolation as prediction, together with linear-scale quantization, Huffman coding, and zstd lossless coding (Collet and Kucherawy, 2018). SZ3 has excellent compression ratio and fidelity over others while its speed may not be as fast as ZFP. SZ3-M compresses the input with different error bounds independently and groups those compressed data all together as output. Such a solution supports multi-fidelity retrieval but is not progressive yet, as it cannot reuse the low-fidelity data to build high-fidelity results.
@@
ZFP-R (Magri and Lindstrom, 2024): ZFP-R is the progressive version of ZFP based on residuals. It handles residuals the same way as SZ3-R. ZFP (Lindstrom, 2014) is the leading transform-based lossy compressor based on orthogonal transformation. ZFP is usually the fastest scientific lossy compressor because of its highly efficient transform function, although its compression ratio may not be as high as others.
@@
We note that the residual-based approaches (SZ3-R and ZFP-R) have drawbacks that affect our evaluation. First, they have limited error-bound flexibility. More specifically, the retrieval is only possible at a few predefined error bounds. This creates a trade-off – setting too many error bounds reduces the overall compression ratio and significantly degrades compression/decompression throughput, while setting too few nodes limits flexibility in selecting error bounds during decompression. For the experiments, we configure five error bounds for them, where each successive one is a factor of $2^{2}$ (4×) apart. Specifically, the error bounds are set to $2^{16}eb$, $2^{14}eb$, $2^{12}eb$, $2^{10}eb$, $2^{8}eb$, $2^{6}eb$, $2^{4}eb$, $2^{2}eb$, and $eb$. Second, those methods are not well-suited for retrievals based on pre-defined bitrates, as the residuals are compressed by error bounds, and their size is not aligned with the bitrate of the retrieval requests. As a result, we select the largest residual that fits within the user-specified bitrate constraint for SZ3-R and ZFP-R in the experiments.